OpenStudy (anonymous):

f(x)=8x√x+3 Find the domain of the function

hartnn (hartnn):

the domain will contain values, for which the quantity under \(\sqrt{....} \) will be positive. so, here, if x+3 >0 what can you say about 'x' ?

OpenStudy (anonymous):

@warsop87 just like the last one, only with \(x+3\geq 0\) instead of \(x-3\geq 0\)

OpenStudy (anonymous):

It told me that was wrong it said the right answer was the infinity sign and 3 so I am confused now

OpenStudy (anonymous):

\[x-3\geq 0\iff x\geq 3\] in interval notation this is \[[3,\infty)\]

OpenStudy (anonymous):

it is just a different notation, interval notation rather than inequality notation

OpenStudy (anonymous):

ohhhh ok

hartnn (hartnn):

for this, \([-3,\infty)\)

OpenStudy (anonymous):

for example \[-2<x\leq 7\] vs \[(-2,7]\]

OpenStudy (anonymous):

thank you so much

hartnn (hartnn):

you are welcome ^_^